# Incentivize Contribution and Learn Parameters Too: Federated Learning with Strategic Data Owners

### v2

@@ 4       Center broadcasts $s^{t}$
6             Perform local training step and compute $[g(w^{0},s^{t},\mu^{t})]_{i}$
@@ 13       Center broadcasts $w^{t}$
15             Compute local gradient: $d_{i}^{t+1}=\nabla_{w}\>a_{i}(w^{t},s_{i}^{*},s_{-i}^{*})$
@@ 16             Send $d_{i}^{t+1}$ to the center
C.2 Model Architecture
Table 1: Layer-wise Summary of CNNModel used for FeMNIST
Layer (type:depth-idx) Output Shape Param #
CNNModel [1, 10]
Conv2d: 1-1 [1, 32, 28, 28] 832
MaxPool2d: 1-2 [1, 32, 14, 14]
Conv2d: 1-3 [1, 64, 14, 14] 51,264
MaxPool2d: 1-4 [1, 64, 7, 7]
Linear: 1-5 [1, 2048] 6,424,576
Linear: 1-6 [1, 10] 20,490
Total Params 6,497,162
Trainable Params 6,497,162
Non-trainable Params 0
Total Mult-Adds (MB) 17.15
Table 2: Layer-wise Summary of CNNModel for CIFAR
Layer (type:depth-idx) Output Shape Param #
CNNModel [1, 10]
Conv2d: 1-1 [1, 32, 32, 32] 896
Conv2d: 1-2 [1, 32, 30, 30] 9,248
MaxPool2d: 1-3 [1, 32, 15, 15]
Conv2d: 1-4 [1, 64, 15, 15] 18,496
Conv2d: 1-5 [1, 64, 13, 13] 36,928
MaxPool2d: 1-6 [1, 64, 6, 6]
Linear: 1-7 [1, 512] 1,180,160
Linear: 1-8 [1, 10] 5,130
Total Params 1,250,858
Trainable Params 1,250,858
Non-trainable Params 0
Total Mult-Adds (MB) 20.83
Input Size (MB) 0.01
Forward/Backward Pass Size (MB) 0.70
Params Size (MB) 5.00
Estimated Total Size (MB) 5.71
Table 3: Layer-wise Summary of Twitter Model
Layer (type:depth-idx) Output Shape Param #
Twitter [1, 2]
Embedding: 1-1 [1, 25, 300] 120,000,300
LSTM: 1-2 [1, 25, 100] 241,600
Linear: 1-3 [1, 128] 12,928
Linear: 1-4 [1, 2] 258
Total Parameters 120,255,086
Trainable Parameters 120,255,086
Non-trainable Params 0
Total Mult-Adds (MB) 126.05
Input Size (MB) 0.00
Forward/Backward Pass Size (MB) 0.08
Params Size (MB) 481.02
Estimated Total Size (MB) 481.10

### v3

@@ 4 Center broadcasts $s^{t}$
6      Perform local training step and compute $g(w^{0},s^{t},\mu^{t})]_{i}$
@@ 13 Center broadcasts $w^{t}$
15      Compute local gradient: $d_{i}^{t+1}=\nabla_{w}\>v_{i}(w^{t},s_{i}^{*},s_{-i}^{*})$
@@ 16      Send $d_{i}^{t+1}$ to the center
The cost function used is linear $c_{i}(s_{i})=c_{i}\cdot s_{i}$, where $c_{i}$ is sampled from $U[0,1]$ for each agent. We choose our valuation function to be $v_{i}(w^{k},s^{k})=r_{i}-\mathcal{L}_{i}(w^{k},s^{k})$ where the $\mathcal{L}_{i}(w^{k},s^{k})$ is the loss evaluated on agents $i$’s test dataset at the $k$-th round of training. The social welfare is given by $\sum_{i\in N}v_{i}(w^{k},s^{k})$. We perform strategy updates using the numerical derivative as follows:
The model parameters obtained by agent $i$ after performing one training step locally on the model initialized with $w^{k-1}$ is denoted by $w_{i}^{k}$. We use the cross-entropy loss function for $L_{i}$} on agent i’s dataset. All the experiments are performed on NVIDIA RTX A5000 and NVIDIA RTX A6000 with a 24GB and 48GB core respectively. The experiments required to obtain the results we report take several days to run. For comparing social welfare against $\beta$ and number of agents, we randomize over 10 runs, sampling an integer $s^{0}_{i}$ from $U[\nicefrac{{s^{\max}_{i}}}{{3}},\nicefrac{{2s^{\max}_{i}}}{{3}}]$ and cost from $U[0,1]$. We plot the average value of time taken across $10$ runs.
For FEMNIST, we use $20\%$ of the dataset that generates users with data that is not identically distributed as per the codebase provided by (Caldas et al., 2019). We train a CNN with $6,603,710$ parameters for $100$ rounds. We choose $\gamma=0.1$ and $\beta=2$. We use the Adam optimizer with learning rate $0.001$.
For the Twitter dataset, we use $15\%$ of the dataset that generates users with data that is identically distributed as per the codebase provided by (Caldas et al., 2019). We further group $1000$ agents together to create agents with larger data shares. We use an LSTM network to train for $50$ rounds. We choose $\gamma=0.5$ and $\beta=2$. We use the Adam optimizer with learning rate $0.00001$.
All results are summarized in Figure˜2.
(a) CIFAR-10
(b) FEMNIST
(c) Twitter
We observe that 2P-UPBReD and FedAvg have the same social welfare performance w.r.t. the number of agents (first column) and running times are nearly the same for all the datasets (third column). 2P-UPBReD outperforms UPBReD in nearly all the runs with respect to social welfare. We also observe that 2P-UPBReD converges to the maximum contribution of the agents for values of $\beta$ that are lower than the theoretically predicted ones. We also note that the data contributors get paid and data consumers pay in 2P-UPBReD and their payments are monotone non-decreasing with data contribution (Figure˜3). The utilities vary because the dataset on which the model is tested to obtain the accuracy varies for the agents.
In Section˜4.2, we describe the trimmed mean algorithm for 2P-UPBReD designed to be robust in the presence of adversarial agents. In this section, we demonstrate this robustness on the FEMNIST and CIFAR-10 dataset. We vary the adversarial fraction $\alpha\in\{0.025,0.05,0.075,0.1,0.125,0.15,0.175,0.2\}$, selecting suitable trimming parameters. Figure˜4 shows the ratio of social welfare with and without adversaries for UPBReD and 2P-UPBReD. While both suffer welfare loss under adversaries, applying the trimmed mean algorithm Algorithm˜4 keeps the welfare ratio close to $1$ for trimmed mean 2P-UPBReD.
(a) CIFAR-10
(b) FEMNIST
Figure 4: Ratio of social welfare with adversaries and without adversaries.
6 Conclusions and future work
In this paper, we proposed UPBReD, which ensures convergence to a Nash equilibrium with model parameters $w$, and its modification T-UPBReD, which enforces truthful reporting. Since $w$ may deviate from the optimal $w^{\texttt{{OPT}}}$, we introduced a two-phase mechanism, 2P-UPBReD, that guarantees full agent participation and convergence to $w^{\texttt{{OPT}}}$, at the cost of monetary transfers that can be internal among agents. Future directions include incorporating agent budget constraints to preserve approximate guarantees and extending the framework to heterogeneous data settings, enabling clustered learning of optimal parameters.